The forces of 10N and 20N are in plain at an angle 60⁰ to each other. Find the resultant and the direction using
1. Analytical method
2. Scle drawing method
1
Expert's answer
2019-10-22T10:19:51-0400

1)


"F_x=F_{1x}+F_{2x}=10+20\cos 60^{\circ}=20N"

"F_y=F_{1y}+F_{2y}=0+20\sin 60^{\circ}=10\sqrt{3}N"

The absolute value of the resultant force


"F=\sqrt{F_x^2+F_y^2}=\sqrt{20^2+(10\sqrt{3})^2}=26.46N"

Direction


"\theta =\tan^{-1}\frac{F_y}{F_x}=\tan^{-1}\frac{10\sqrt{3}}{20}=41^{\circ}"

2)